Grade 11General Physics

A monkey of mass 20 Kg is holding a vertical rope. The rope can break when a mass of 25 Kg is suspended

from it. What is the maximum acceleration with which the monkey can climb up along the rope?

(a)2.5 ms-2 (b) 5 ms-2 (c) 7 ms-2 (d) 10 ms-2

jeet jeetu
13 Years agoGrade 11

8 Answers

Anirban Mukherjee
ApprovedApproved Tutor Answer13 Years ago
Let the monkey climb up with acceleration a. Then tension caused in the rope will be T =

M(g+a), where M is the mass of the monkey. Maximum value of T is 25g = 20(g + a). Taking

g = 10ms

-2, it gives a = 2.5 ms
Aman  Bansal
13 Years ago

Dear Jeet,

assuming the value of g is 10ms^-2, the maximum weight that the rope can hold is  

25kg×10Nkg1=250N

. (though it should be less than 250, I''m not so sure about the maimum part) using F=ma,

a=250N20kg =12.5ms2

The maximum acceleration of the monkey is 12.5ms^-2

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert

somy chaturvedi
13 Years ago

tension is responsiblefor upward acc. and  maximum tension it can bear is 250N. hence max acc.

250-200=20*a

a=2.5m/s^2

sheik arbaz
13 Years ago

2.5 m/s-2

Shahvar Tyagi
9 Years ago
A monkey of mass 20 kg. is holding a vertical rope. The rope will not break when a mass of 25 kg. is suspended from it but will break if the mass exceeds 25 kg. What is the maximum acceleration it which the monkey can climb up along the rope
Ajinkya kawale
9 Years ago
let the initial tension be Tusing fbd for m=20kg we get,T-mg=maT-20g=20a…1Tmax-25g=0Tmax=25g……2from 1,225g-20g=20a5g=20aa=g/4 =2.5m/s^2hope this helps…
Vicky Singh
7 Years ago
1kgf = 1kg * 10m/s square 
         =10N
For 25kgf=25kg * 10m/s square 
                =250N _____(1)
 Monkey mass =20kg so mg=200N (2)
      Subtract equation 2 from 1 we get 
        250N-200N=50N
      Force = Mass *acceleration 
       50 =20kg * a
       a = 2.5 m/s square 
              
Rishi Sharma
5 Years ago
Dear Student,
Please find below the solution to your problem.

1kgf = 1kg * 10m/s square
=10N
For 25kgf=25kg * 10m/s square
=250N _____(1)
Monkey mass =20kg so mg=200N (2)
Subtract equation 2 from 1 we get
250N-200N=50N
Force = Mass *acceleration
50 =20kg * a
a = 2.5 m/s square

Thanks and Regards